40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Feb 2 18:35:47 2016, gautam, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table 
    Reply  Tue Feb 2 18:54:02 2016, Koji, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table 
       Reply  Wed Feb 3 23:23:47 2016, gautam, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table KnifeEdgeScans_x.pdfBeamscan_x.pdf
          Reply  Thu Feb 4 09:16:46 2016, Koji, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table 
             Reply  Fri Feb 5 00:23:01 2016, gautam, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table Beamscan_setup.pdfBeamscan_x.pdfBeamscan_y.pdfZscan.pdf
                Reply  Fri Feb 5 16:50:24 2016, gautam, Update, Green Locking, First pass at mode-matching Modematch_AUXx.pdfNewSetUp.pngModematch_AUXx_2.pdfXendModeMatch.m.zip
                   Reply  Mon Feb 8 15:36:37 2016, gautam, Update, Green Locking, Alternative mode-matching scheme Modematch_alternative.pdf
                      Reply  Tue Feb 9 04:37:10 2016, ericq, Update, Green Locking, Laser swap initiated 
                         Reply  Tue Feb 9 11:49:47 2016, gautam, Update, Green Locking, Laser swap initiated Innolight_beamProfile.pdf
                            Reply  Tue Feb 9 19:15:36 2016, gautam, Update, Green Locking, Laser swap - updates Modematch_X.pdfXendModeMatch.m.zip
                               Reply  Wed Feb 10 17:57:15 2016, gautam, Update, Green Locking, Laser swap - updates 
                                  Reply  Fri Feb 12 17:05:40 2016, gautam, Update, Green Locking, Laser swap - green light recovered but no flashes in the arm 
                                     Reply  Fri Feb 12 19:07:52 2016, Koji, Update, Green Locking, Laser swap - green light recovered but no flashes in the arm 
                                        Reply  Fri Feb 19 00:42:19 2016, gautam, Update, Green Locking, Laser swap - beam ellipticity from laser? IMG_6287.JPGIMG_6288.JPGIMG_6286.JPG
                                           Reply  Mon Feb 22 13:56:52 2016, gautam, Update, Green Locking, Laser swap -reflected beam from ETM aligned 
                                              Reply  Tue Feb 23 23:01:16 2016, gautam, Update, Green Locking, Laser swap - Green PDH locking AM_scan.pdfAM_scan_zoomed.pdf
                                                 Reply  Wed Feb 24 19:29:13 2016, gautam, Update, Green Locking, Laser swap - Green PDH locked 
                                                    Reply  Fri Feb 26 01:52:44 2016, gautam, Update, Green Locking, Laser swap - Green PDH OLTF X_PDH_OLTF_20160225.pdf
                                                       Reply  Mon Feb 29 17:17:26 2016, gautam, Update, Green Locking, Laser swap - still no green beatnote 
                                                          Reply  Wed Mar 2 17:42:19 2016, gautam, Update, Green Locking, Laser swap - some progress 
                                                             Reply  Thu Mar 3 01:25:50 2016, gautam, Update, Green Locking, Laser swap - 2 IR + 1 green beatnotes found IR_beat_20160303_2.pdf
                                                                Reply  Fri Mar 4 01:11:41 2016, gautam, Update, Green Locking, Laser swap - both green beatnotes found IR_beat_20160303_green.pdf
                                                                   Reply  Sat Mar 5 23:31:01 2016, gautam, Update, Green Locking, Laser swap - some updates AUXxTelescope.png.pngPM_setup.pdfPLLolg.pdfAMPM20160303.pdfIRbeat_20160304.pdf
                                                                      Reply  Mon Mar 7 23:51:36 2016, gautam, Update, Green Locking, Laser swap - some improvement PomonaTF.pdfXPDH.pdfYPDH.pdfgreenbeat_20160307.pdf
                                                                         Reply  Tue Mar 8 18:22:20 2016, rana, Update, Green Locking, Laser swap - some improvement 
Message ID: 11999     Entry time: Fri Feb 19 00:42:19 2016     In reply to: 11989     Reply to this: 12002
Author: gautam 
Type: Update 
Category: Green Locking 
Subject: Laser swap - beam ellipticity from laser? 

Eric and I spent some time yesterday night trying to recover the green in the arm after the laser swap. The problem essentially was that though I was getting ~800uW of green out of the doubling oven, the mode wasn't clean, and hence, the beam profile looked really messed up just before entering the arm cavity.We got to a point where we thought we were getting a good mode out of the doubling oven (as judged by propagating this beam onto the wall with the help of a mirror). But we were only getting ~400uW of green power. I tried tweaking the alignment of the oven on the 4 axis stage for a while, but was not able to improve the situation much. So I decided to start from scratch:

  • First, I made sure that the IR beam from the laser was hitting the first steering mirror approximately at the center (see here for the optical layout). 
  • Then I used the two steering mirrors immediately after the laser to make sure that the IR beam was hitting the first lens and the HWP before the IR faraday roughly at their center. 
  • Next, I propagated the beam through the IR Faraday, again using SM1 and SM2 to do the steering - initial alignment through the Faraday was done by eye, and I did  some fine adjustment by maximizing the power coming out of the Faraday. We have 252mW of IR going into the IR Faraday, and 225mW coming out. I judged that these numbers were reasonable, and compared favourably to what we had with the Innolight.
  • Keeping the downstream alignment, I used SM1 and SM2 to hit the second lens roughly at its center. I then re-measured the distance from this lens to the center of the doubling oven, and tweaked this slightly to match my mode-matching calculation. 
  • I then tried to carefully play with this lens and the alignment of the doubling oven using the four axis stage. After (many) iterations, and with some luck, I managed to find what I judged to be a good alignment. Using the mirror-on-a-stick to reflect the green beam out of the doubler onto the wall nearby (see Attachment #1, all photos taken using my phone camera), the mode looks reasonably clean. I was also able to get 1mW of green power out of the doubler, an efficiency of ~2%/W. The doorknob should give some sense of scale, but at this point, the mode looks pretty clean (this was not the case previously).
  • I then aligned the post doubler optics to send the beam through the green Faraday (~0.85mW of green out of the green Faraday) and through the two irides I had put in before swapping the lasers. As the beam propagates, however, some ellipticity in it becomes more and more apparent - especially after the f=-100mm lens between the two piezo mirrors. Attachment #2 shows the beam immediately after this lens, while Attachment #3 shows the beam on the iris just before it is sent into the arm cavity. 

I am beginning to wonder if this ellipticity is inherent from the IR beam from the laser? My beamscan results suggest that the beam is more divergent in the "P direction" as compared to the "S direction", which is borne out by these photographs. And if this is indeed the case, do we need to add cylindrical lenses to correct this?


Unrelated to this work: The ITMX Oplev seems to have wandered off so the X arm won't lock. I am not realigning the Oplev for now, but am turning the ITMX Oplev servo off for the night. 

Perhaps related to my work on the endtable: The ETMX oplev MEDM readings seemed to be frozen, though there was red light on the QPD on the endtable. Checking the CDS overview screen, I saw that all models on c1iscex had crashed. I sshed into c1iscex and restarted all the models, but the IOP block remained red. I checked the datetime, and found that this was wrong - so I followed the instructions here, but the "Diag Word" block remains red. I am shutting down the watchdog for ETMX and leaving this as is for now... This seems to have happened before...

Attachment 1: IMG_6287.JPG  238 kB  | Hide | Hide all | Show all
IMG_6287.JPG
Attachment 2: IMG_6288.JPG  315 kB  | Hide | Hide all | Show all
IMG_6288.JPG
Attachment 3: IMG_6286.JPG  294 kB  | Show | Hide all | Show all
ELOG V3.1.3-